# SPECIAL RELATIVITY

### v9

@@ 14.2.3 The Motion of an Electron Accelerated by a Kicker in a Bending Magnet
In our relativistic but non-covariant study of electron motion in a given magnetic field, the electron has the same velocity and consequently the same relativistic factor $\gamma$ upstream and downstream of the kicker. Suppose we now put the electron through a bending magnet (i.e. a uniform magnetic field directed along the $y$-axis ), Fig. 62. The motion in the bending magnet we obtained is practically the same as in the case of non-relativistic dynamics, the only difference being the appearance of the relativistic factor $\gamma$ in the determination of cyclotron frequency $\omega_{c}=eB/(m\gamma)$. The curvature radius $R$ of the trajectory is derived from the relation $v_{\perp}/R=\omega_{c}$, where $v_{\perp}=v(1-\theta_{k}^{2}/2)$ is the component of the velocity normal to the field of the bending magnet $\vec{B}=B\vec{e}_{y}$. According to non-covariant particle tracking, after the kick, the correction to the radius $R$ is only of order $\theta_{k}^{2}$.
@@
Figure 62: Geometry for radiation production from a bending magnet. The motion of a relativistic electron accelerated by a kicker field. According to the non-covariant particle tracking , the magnetic field $B\vec{e}_{x}$ of the kicker is only capable altering the direction of motion, but not the speed of the electron.
@@
One could naively expect that according to covariant particle tracking, since the total speed of electron in the lab frame downstream of the kicker decreases from $v$ to $v(1-\theta_{k}^{2}/2)$, this would also lead to a consequent decrease of the three-momentum $|\vec{p}|$ from $m\gamma v$ to $m\gamma v(1-\gamma^{2}\theta_{k}^{2}/2)$ in our approximation. However, such a momentum change would mean a correction to the radius $R$ of order $\gamma^{2}\theta_{k}^{2}$ so that there is a glaring conflict with the calculation of the radius according to non covariant tracking. Since the curvature radius of the trajectory in the bending magnet has obviously an objective meaning, i.e. it is convention-invariant, this situation seems paradoxical. The paradox is solved taking into account the fact that in Lorentz coordinates the three-vector of momentum $\vec{p}$ is transformed, under Lorentz boosts, as the space part of the four vector $p_{\mu}$. Let us consider a composition of Lorentz boosts that track the motion of the relativistic electron accelerated by the kicker field. Under this composition of boosts the longitudinal momentum component remains unchanged (with accuracy $\theta_{k}^{2}$).
@@
Let us verify that this assertion is correct. We have $p_{\mu}=[\mathcal{E}/c,\vec{p}]$. We consider the Lorentz frame $S^{\prime}$ fixed with respect to the electron upstream the kicker, and in the special case when electron is at rest $p_{\mu}^{\prime}=[mc,\vec{0}]$. We turn focus on what happens in $S^{\prime}$. Acceleration in the crossed kicker fields gives rise to an electron velocity $v_{y}^{\prime}=\gamma v_{y}$ parallel to the $y$-axis and $v_{z}^{\prime}=-v(\gamma v_{y})^{2}/2$ parallel to the $z$-axis. Downstream of the kicker the transformed four-momentum is $p_{\mu}^{\prime}=[mc+mv_{y}^{\prime 2}/(2c),0,mv_{y}^{\prime},mv_{z}^{\prime}]$, where we evaluate the transformation only up to the order $(\gamma v_{y}/c)^{2}$, as done above. We note that, due to the transverse boost, there is a contribution to the time-like part of the four-momentum vector i.e. to the energy of the electron. In fact, the energy increases from $mc^{2}$ to $mc^{2}+m(\gamma v_{y})^{2}/2$. We remind that $S^{\prime}$ is connected to the lab frame $S$ by a Lorentz boost. Now, with a boost to a frame moving at velocity $\vec{v}=-v\vec{e}_{z}$, the transformation of the longitudinal momentum component, normal to the magnetic field of the bend, is $p_{z}=\gamma(p_{z}^{\prime}+vp_{0}^{\prime}/c)=\gamma mv$. Therefore we can see that the momentum component along the $z$-axis remains unchanged in our approximation of the Lorentz transformation. We also have, from the transformation properties of four-vectors, that the time component $p_{0}=\gamma(p^{\prime}_{0}+vp^{\prime}_{z})=\gamma mc$
@@
Let us now return to our consideration on the covariant electron trajectory calculation in the Lorentz lab frame when a constant magnetic field is applied. We analyzed a very simple (but very practical) kicker setup and we noticed that, in fact, the three-momentum is not changed; so we have already verified that this transformation is the same as the non covariant transformation for the three-momentum, i.e. $\vec{p}_{cov}=\vec{p}$. We also found that there is a difference between covariant and non covariant output velocities, $v_{cov}<v$. In these transformations we therefore demonstrated that $\vec{p}_{cov}\neq m\vec{v}_{cov}/\sqrt{1-v^{2}_{cov}/c^{2}}$ for curved trajectory in ultrarelativistic asymptotic. It is interesting to discuss what it means that there are two different (covariant and non covariant) approaches that produce the same particle three-momentum. The point is that both approaches describe correctly the same physical reality and the curvature radius of the trajectory in the magnetic field (and consequently the three-momentum) has obviously an objective meaning, i.e. is convention-invariant. In contrast to this, the velocity of the particle has objective meaning only up to a certain accuracy, because the finiteness of velocity of light takes place.
14.2.4 Red Shift of the Critical Frequency of the Synchrotron Radiation
@@
Next we discuss the interesting problem of emission of synchrotron radiation in a bending magnet with and without kick. Let us consider the setup pictured in Fig. 62. Suppose that an ultrarelativistic electron moving along the $z$-axis in the lab frame is kicked by a weak dipole field directed along the $x$-axis before entering a uniform magnetic field directed along the $y$-axis, i.e. a bending magnet. An accelerated electron traveling on a curved trajectory emits radiation. When moving at relativistic speed, this radiation is emitted as a narrow cone tangent to the path of the electron. Moreover, the radiation amplitude becomes very large in this direction. This phenomenon is known as Doppler boosting. Synchrotron radiation is generated when a relativistic electron is accelerated in a bending magnet. Without going into the details of computation, it is possible to present intuitive arguments explaining why the characteristics of the spectrum of synchrotron radiation only depend, in the ultrarelativistic limit, on the difference between electron and light speed.
@@
We turn now to the radiation emitted by an ultrarelativistic electron in a bending magnet. Let us discuss the case when the source is heading towards the observer (Fig. 63). An electromagnetic source propagates through the system as a function of time following a certain trajectory $\vec{x}(t^{\prime})$. However, an electromagnetic signal emitted at time $t^{\prime}$ at a given position $\vec{x}(t^{\prime})$ arrives at the observer position at a different time $t$, due to the finite speed of light. As a result, an observer sees the motion of the electromagnetic source as a function of $t$. The prime used here to indicate the retarded times should not be confused with the primes referring to a Lorentz transformed frame in the proceeding sections.
@@
Figure 63: Geometry for synchrotron radiation from a bending magnet. Radiation from an electron passing through the setup is observed through a spectral filter by a fixed observer positioned on the tangent to the bend at point $P$, as shown in Fig. (a). Electromagnetic source propagates through the system, as a function of time, as shown in Fig. (b). However, electromagnetic signal emitted at time $t^{\prime}$ at a given position $x(t^{\prime})$ arrives at observer position at a different time $t$, due to finite speed of light. As a result, the observer in Fig. (a) sees the electromagnetic source motion as a function of $t$. The apparent motion is a hypocycloid, and not the real motion $x(t^{\prime})$. The observer sees a time-compressed motion of the sources, which go from point $A$ to point $B$ in an apparent time corresponding to an apparent distance $2R\theta dt/dt^{\prime}$.
Figure 64: Formation length for bend. Formation length $L_{f}$ can also be considered as a longitudinal size of a single-electron source. It does not make sense at all to talk about the position where electromagnetic signals are emitted within $L_{f}$.
@@
It should be note, however, that there is another satisfactory way of explaining the red shift. We can reinterpret this result with the help of a non-covariant treatment, which deals with non- covariant particle trajectories, and with Galilean transformations of the electromagnetic field equations. According to non-covariant particle tracking the electron velocity is unvaried. However, Maxwell’s equations do not remain invariant with respect to Galilean transformation, and the velocity of light has increased from $c$, without kick, to $c(1+\theta_{k}^{2}/2)$ with kick. The reason for the velocity of light being different from the electrodynamics constant $c$ is due to the fact that, according to the absolute time convention, the clocks after the kick are not resynchronized. The ratio of the electron velocity to that of light is convention independent i.e. it does not depend on the distant clock synchronization or on the rhythm of the clocks. Our calculations show that covariant and non-covariant treatments (at the correct coupling fields and particles) give the same result for the red shift prediction, which is obviously convention-invariant and depends only on the (dimensionless) parameter $v_{coord}/c_{coord}=v(1-\theta^{2}_{k}/2)/c=v/[c(1+\theta^{2}_{k}/2)]$, where $v_{coord}$ is the coordinate electron velocity, and $c_{coord}$ is the coordinate velocity of light.
In order to confirm the predictions of our synchrotron radiation theory, we propose an experimental test at third generation synchrotron radiation sources. Synchrotron radiation from bending magnets is emitted within a wide range of frequencies. The possibility of using narrow bandwidth sources in an experimental study on the red shift in synchrotron radiation spectrum looks more attractive. This allows one to increase the sensitivity of the output intensity on the red shift, and to relax the requirement on beam kicker strength and photon beam line aperture. Undulators, as sources of quasi-monochromatic synchrotron radiation, produce light in a sufficiently narrow bandwidth for our purposes. They cause the electron beam to follow a periodic undulating trajectory with the consequence that interference effects occur. Undulators have typically many periods. The interference of radiation produced in different periods results in a bandwidth that scales as the inverse number of periods. Therefore, the use of insertion devices installed at third generation synchrotron radiation facilities would allow us to realize a straightforward increase in the sensitivity to the red shift at a relatively small kick angle, $\theta_{k}<1/\gamma$. In the next chapter we discuss this experimental test in more detail.
14.3 Bibliography and Notes
@@ 15.1 Introductory Remarks
Accelerator physics was always thought in terms of the old (Newtonian) kinematics that is not compatible with Maxwell’s equations. We would like now to use our ideas about dynamics and electrodynamics to consider in some greater detail the question: ”Why did the error in the synchrotron radiation theory remain so long undetected?”
@@
The phenomena of electromagnetic radiation which we want to study are relatively complicated. For an arbitrary setup covariant calculations of the radiation process is very difficult. There are, however, circumstances in which calculations can be greatly simplified. As example of such circumstance is a synchrotron radiation setup. Similar to the non-relativistic asymptote, the ultrarelativistic asymptote also provides the essential simplicity of the covariant calculation. The reason is that the ultrarelativistic assumption implies the paraxial approximation. Since the formation length of the radiation is much longer than the wavelength, the radiation is emitted at small angles of order $1/\gamma$ or even smaller, and we can therefore enforce the small angle approximation. We assume that the transverse velocity is small compared to the velocity of light. In other words, we use a second order relativistic approximation for the transverse motion. Instead of small (total) velocity parameter $(v/c)$ in the non-relativistic case, we use a small transverse velocity parameter $(v_{\perp}/c)$. The next step is to analyze the longitudinal motion, following the same method. We should remark that the analysis of the longitudinal motion in a synchrotron radiation setup is very simple. If we evaluate the transformations up to second order $(v_{\perp}/c)^{2}$, the relativistic correction in the longitudinal motion does not appear in this approximation.
@@
According to covariant approach, the various relativistic kinematics effects concerning to the synchrotron radiation setup, turn up in successive orders of approximation.
In the first order $(v_{\perp}/c)$. - relativity of simultaneity. Wigner rotation, which in the ultrarelativistic approximation appears in the first order already, and results directly from the relativity of simultaneity.
In the second order $(v_{\perp}/c)^{2}$. - time dilation. Relativistic correction in law of composition of velocities, which already appears in the second order, and results directly from the time dilation.
@@
The first order kinematics term $(v_{\perp}/c)$ plays an essential role only in the description of the coherent radiation from a modulated electron beam. In a storage ring the distribution of the longitudinal position of the electrons in a bunch is essentially uncorrelated. In this case, the radiated fields due to different electrons are also uncorrelated and the average power radiated is a simple sum of the radiated power from individual electrons; that is we sum intensities, not fields. In the case of incoherent synchrotron radiation, a motion of the single ultrarelativistic electron in a constant magnetic field, according to the theory of relativity, influences the kinematics terms of the second order $(v_{\perp}/c)^{2}$ only.

### v10

@@ 18.3 The Motion of an Electron Accelerated by a Kicker in a Bending Magnet
In our relativistic but non-covariant study of electron motion in a given magnetic field, the electron has the same velocity and consequently the same relativistic factor $\gamma$ upstream and downstream of the kicker. Suppose we put the electron through a bending magnet (i.e. a uniform magnetic field directed along the $y$-axis ), Fig. 51. The motion in the bending magnet we obtained is practically the same as in the case of non-relativistic dynamics, the only difference being the appearance of the relativistic factor $\gamma$ in the determination of cyclotron frequency $\omega_{c}=eB/(m\gamma)$. The curvature radius $R$ of the trajectory is derived from the relation $v_{\perp}/R=\omega_{c}$, where $v_{\perp}=v(1-\theta_{k}^{2}/2)$ is the component of the velocity normal to the field of the bending magnet $\vec{B}=B\vec{e}_{y}$. According to non-covariant particle tracking, after the kick, the correction to the radius $R$ is only of order $\theta_{k}^{2}$.
@@
Figure 51: Geometry for radiation production from a bending magnet. According to the non-covariant particle tracking, the magnetic field $B\vec{e}_{x}$ of the kicker is only capable altering the direction of motion, but not the speed of the electron.
@@
One could naively expect that according to covariant particle tracking since the total speed of an electron in the lab frame downstream of the kicker decreases from $v$ to $v(1-\theta_{k}^{2}/2)$, this would also lead to a consequent decrease of the three-momentum $|\vec{p}|$ from $m\gamma v$ to $m\gamma v(1-\gamma^{2}\theta_{k}^{2}/2)$ in our approximation. However, such a momentum change would mean a correction to the radius $R$ of order $\gamma^{2}\theta_{k}^{2}$ so that there is a glaring conflict with the radius calculation according to noncovariant tracking. Since the curvature radius of the trajectory in the bending magnet has an objective meaning, i.e. it is convention-invariant, this situation seems paradoxical. The paradox is solved taking into account the fact that in Lorentz coordinates the three-vector of momentum $\vec{p}$ is transformed, under Lorentz boosts, as the space part of the four-vector $p_{\mu}$. Let us consider a composition of Lorentz boosts that track the motion of the relativistic electron accelerated by the kicker field. Under this composition of boosts the longitudinal momentum component remains unchanged (with accuracy $\theta_{k}^{2}$).
@@
Let us verify that this assertion is correct. We have $p_{\mu}=[\mathcal{E}/c,\vec{p}]$. We consider the Lorentz frame $S^{\prime}$ fixed with respect to the electron upstream the kicker, and in the special case when an electron is at rest $p_{\mu}^{\prime}=[mc,\vec{0}]$. We turn focus on what happens in $S^{\prime}$. Acceleration in the crossed kicker fields gives rise to an electron velocity $v_{y}^{\prime}=\gamma v_{y}$ parallel to the $y$-axis and $v_{z}^{\prime}=-v(\gamma v_{y})^{2}/2$ parallel to the $z$-axis. Downstream of the kicker the transformed four-momentum is $p_{\mu}^{\prime}=[mc+mv_{y}^{\prime 2}/(2c),0,mv_{y}^{\prime},mv_{z}^{\prime}]$, where we evaluate the transformation only up to the order $(\gamma v_{y}/c)^{2}$, as done above. We note that, due to the transverse boost, there is a contribution to the time-like part of the four-momentum vector i.e. to the energy of the electron. The energy increases from $mc^{2}$ to $mc^{2}+m(\gamma v_{y})^{2}/2$. We remember that $S^{\prime}$ is connected to the lab frame $S$ by a Lorentz boost. Now, with a boost to a frame moving at velocity $\vec{v}=-v\vec{e}_{z}$, the transformation of the longitudinal momentum component, normal to the magnetic field of the bend, is $p_{z}=\gamma(p_{z}^{\prime}+vp_{0}^{\prime}/c)=\gamma mv$. Therefore we can see that the momentum component along the $z$-axis remains unchanged in our approximation of the Lorentz transformation. We also have, from the transformation properties of four-vectors, that the time component $p_{0}=\gamma(p^{\prime}_{0}+vp^{\prime}_{z})=\gamma mc$.
@@
It is interesting to discuss what it means that there are two different (covariant and noncovariant) approaches that produce the same particle three-momentum. The point is that both approaches describe correctly the same physical reality and the curvature radius of the trajectory in the magnetic field (and consequently the three-momentum) has an objective meaning, i.e. is convention-invariant. In contrast to this, the velocity of the particle has objective meaning only up to a certain accuracy, because the finiteness of the velocity of light takes place.
18.4 Redshift of the Critical Frequency of the Synchrotron Radiation
@@
Next, we discuss the interesting problem of the emission of synchrotron radiation in a bending magnet with and without the kick (see the next chapter). Let us consider the setup pictured in Fig. 51. Suppose that an ultrarelativistic electron moving along the $z$-axis in the lab frame is kicked by a weak dipole field directed along the $x$-axis before entering a uniform magnetic field directed along the $y$-axis, i.e. a bending magnet. An accelerated electron traveling on a curved trajectory emits radiation. When moving at relativistic speed, this radiation is emitted as a narrow cone tangent to the path of the electron. Moreover, the radiation amplitude becomes very large in this direction. This phenomenon is known as Doppler boosting. Synchrotron radiation is generated when a relativistic electron is accelerated in a bending magnet. Without going into the details of computation, it is possible to present intuitive arguments explaining why the characteristics of the spectrum of synchrotron radiation only depend, in the ultrarelativistic limit, on the difference between electron and light speed.
@@
We turn now to the radiation emitted by an ultrarelativistic electron in a bending magnet. Let us discuss the case when the source is heading towards the observer (Fig. 52). An electromagnetic source propagates through the system as a function of time following a certain trajectory $\vec{x}(t^{\prime})$. However, an electromagnetic signal emitted at time $t^{\prime}$ at a given position $\vec{x}(t^{\prime})$ arrives at the observer position at a different time $t$, due to the finite speed of light. As a result, an observer sees the motion of the electromagnetic source as a function of $t$. The prime used here to indicate the retarded times should not be confused with the primes referring to a Lorentz-transformed frame in the proceeding sections.
@@
Figure 52: Geometry for synchrotron radiation from a bending magnet.
Figure 53: Formation length for the bend.
@@
It should be noted, however, that there is another satisfactory way of explaining the redshift. We can reinterpret this result with the help of a non-covariant treatment, which deals with non-covariant particle trajectories, and with Galilean transformations of the electromagnetic field equations. According to non-covariant particle tracking the electron velocity is unvaried. However, Maxwell’s equations do not remain invariant with respect to Galilean transformation, and the velocity of light has increased from $c$, without kick, to $c(1+\theta_{k}^{2}/2)$ with kick. The reason for the velocity of light being different from the electrodynamics constant $c$ is because, according to the absolute time convention, the clocks after the kick are not resynchronized. The ratio of the electron velocity to that of light is convention independent i.e. it does not depend on the distant clock synchronization or the rhythm of the clocks. Our calculations show that covariant and non-covariant treatments (at the correct coupling fields and particles) give the same result for the redshift prediction, which is convention-invariant and depends only on the (dimensionless) parameter $v_{coord}/c_{coord}=v(1-\theta^{2}_{k}/2)/c=v/[c(1+\theta^{2}_{k}/2)]$, where $v_{coord}$ is the coordinate electron velocity, and $c_{coord}$ is the coordinate velocity of light.
To confirm the predictions of our synchrotron radiation theory, we propose an experimental test at third-generation synchrotron radiation sources. Synchrotron radiation from bending magnets is emitted within a wide range of frequencies. The possibility of using narrow bandwidth sources in an experimental study on the redshift in the synchrotron radiation spectrum looks more attractive. This allows one to increase the sensitivity of the output intensity on the redshift, and to relax the requirement on beam kicker strength and photon beam line aperture. Undulators, as sources of quasi-monochromatic synchrotron radiation, produce light in a sufficiently narrow bandwidth for our purposes. They cause the electron beam to follow a periodic undulating trajectory with the consequence that interference effects occur. Undulators typically have many periods. The interference of radiation produced in different periods results in a bandwidth that scales as the inverse number of periods. Therefore, the use of insertion devices installed at third-generation synchrotron radiation facilities would allow us to realize a straightforward increase in the sensitivity to the redshift at a relatively small kick angle, $\theta_{k}<1/\gamma$. In other words, we are stating that, despite the many measurements done over decades, synchrotron radiation theory is not an experimentally well-confirmed theory.
18.5 Discussion
@@
Let us quickly look at some other cases of interest. Above we consider the simple case when the kick angle is small. When the electron passes through the weak kicker there is no synchrotron radiation (to be more precise, in this case, radiation is indistinguishable from the self-electromagnetic fields of the electron). We expect that an electron that passes through the weak kicker is still ”field-dressed”. Suppose the ultrarelativistic electron in the lab frame is kicked by a dipole magnet field, but the kick angle is large compared with $1/\gamma$. In other words, we now consider an electron moving along a standard synchrotron radiation setup. At the exit of the strong kicker, we have a ”naked” (or ”field-free”) electron i.e. an electron that is not accompanied by virtual radiation fields. There is a process of formation of the ”field dressed” electron (i.e. formation of the fields from a fast-moving charge) within the formation length from the very beginning of the straight section downstream of the strong kicker. The information about acceleration is included now in the synchrotron radiation. But the acceleration history (together with the self fields of the ultrarelativistic electron) is washed out during the kicking process, and now electron fields not include information about acceleration. That is to say, the information about acceleration do not included in the covariant trajectory downstream of the strong kicker (we will discuss this subject in the next chapter).
@@ 19.1 Introductory Remarks
Accelerator physics was always thought of in terms of the old (Newtonian) kinematics that is not compatible with Maxwell’s equations. We would like now to use our ideas about dynamics and electrodynamics to consider in some greater detail the question: ”Why did the error in the synchrotron radiation theory remain so long undetected?”
@@
The phenomena of electromagnetic radiation which we want to study are relatively complicated. For an arbitrary setup, covariant calculations of the radiation process are very difficult. There are, however, circumstances in which calculations can be greatly simplified. An example of such a circumstance is a synchrotron radiation setup. Similar to the non-relativistic asymptote, the ultrarelativistic asymptote also provides the essential simplicity of the covariant calculation. The reason is that the ultrarelativistic assumption implies the paraxial approximation. Since the formation length of the radiation is much longer than the wavelength, the radiation is emitted at small angles of order $1/\gamma$ or even smaller, and we can therefore enforce the small angle approximation. We assume that the transverse velocity is small compared to the velocity of light. In other words, we use a second-order relativistic approximation for the transverse motion. Instead of a small (total) velocity parameter $(v/c)$ in the non-relativistic case, we use a small transverse velocity parameter $(v_{\perp}/c)$. The next step is to analyze the longitudinal motion, following the same method. We should note that the analysis of the longitudinal motion in a synchrotron radiation setup is very simple. If we evaluate the transformations up to second order $(v_{\perp}/c)^{2}$, the relativistic correction in the longitudinal motion does not appear in this approximation.
@@
According to the covariant approach, the various relativistic kinematics effects concerning to the synchrotron radiation setup, turn up in successive orders of approximation. In the first order $(v_{\perp}/c)$. - relativity of simultaneity. Wigner rotation, which in the ultrarelativistic approximation appears in the first order already, and results directly from the relativity of simultaneity. In the second order $(v_{\perp}/c)^{2}$. - time dilation. Relativistic correction in the law of composition of velocities, which already appears in the second order, results directly from the time dilation.
@@
The first-order kinematics term $(v_{\perp}/c)$ plays an essential role only in the description of the coherent radiation from a modulated electron beam. In a storage ring, the distribution of the longitudinal position of the electrons in a bunch is essentially uncorrelated. In this case, the radiated fields due to different electrons are also uncorrelated and the average power radiated is a simple sum of the radiated power from individual electrons; that is we sum intensities, not fields. In the case of incoherent synchrotron radiation, a motion of the single ultrarelativistic electron in a constant magnetic field, according to the theory of relativity, influences the kinematics terms of the second order $(v_{\perp}/c)^{2}$ only.